Almost all TPC sectors tripped, except of VTPC-1 sectors 4 and 5, and some sectors in MTPC-R. BPDs tripped also.
-> Power supply problem (it trips even if the cable is disconnected). Rainer is arriving.
20:30 Solution is found. Accidentally the current limit for all channels in the CAEN power supply was set to 10nA. Now the original settings are restored and we could turn on all TPCs, BPDs, counters, potential strips (they tripped also). |